Answer all questions !
9-5: Constant Growth Stocks Valuation of a constant growth stock A stock is expected to pay a dividend of $1.75 the end of the vear that is. D.-$1.75 and it should continue t grow at a constant rate of 7% a year. If ts re are returns 13%, what is the stocks expected pnce 4 years from today 7 Round your answer to two decimal places.
media%2F89a%2F89a4a9a6-ba74-4018-a6cb-d6
media%2F307%2F307ca3f7-8881-4336-a290-1c
Answer #1

Question 1

D1 = 1.75

D2 = 1.75 * (1 + 7%) = 1.8725

D3 = 1.8725 * (1 + 7%) = 2.0036

D4 = 2.0036 * (1 + 7%) = 2.1438

D5 = 2.1438 * (1 + 7%) = 2.2939

5

V4 = 2.2939/(13% - 7%) = $38.23--> Answer

Question 2

Based on CAPM,

Cost of Equity = Risk free rate + Beta * Market risk premium

Cost of equity (r) = 5.3% + (0.9 * 4.5%) = 9.35%

Based on constant growth dividend discount model,

V_0 = \frac{D_1}{r - g}

48 = 2.50/(9.35% - g)

9.35% - g = 5.21%

g = 4.14%

D4 = D1 * (1 + g)3

D4 = 2.50 * (1 + 4.14%)3 = $2.8237

V3 = D4/(r - g) = 2.8237/(9.35% - 4.14%) = $54.21

Question 3

Based on constant growth dividend discount model,

V_0 = \frac{D_1}{r - g}

D1 = 1.25 * (1 + 6%) = $1.325

34 = 1.325/(r - 6%)

r = 3.90% + 6% = 9.90% --> Answer for part b

2 T- g

D2 = 1.325 * (1 + 6%) = 1.4045

V1 = 1.4045/(9.90% - 6%) = $36.01 --> Answer to part a